A block slides down an incline of angle 30o with an acceleration of g/4. find the kinetic friction coefficient

Grade:10

3 Answers

kedar joshi
13 years ago

let u be the coefficient of kinetic friction.

(netforce)mg/4=mgsin@-umgcos@

by solving we get

sin@-ucos@=1/4

by substituting @=30 we get

u=2/root of 3

please approve if you feel the answer is correct

Suchismita Paul
6 years ago
Answer should be u=1/2root3 because mg(sin@-ucos@)=mg/4sin 30 - ucos 30=1/41/2-uroot3/2=1/4u = 1/2root3
Ritesh Ranjan
6 years ago
Given :Inclination angle=θ=30 degreesa=g/4 m/s²Coefficient of Kinetic friction .Total net force= mg/4=mgsinθ-μmgcosθwhere μ is coefficient of kinetic friction.mg/4=mg[sinθ-μcosθ][sinθ-μcosθ] =1/4Substituting the inclination angle, we getsin30- μcos30=1/41/2-μ√3/2=1/41/2-1/4=μ√3/22-1/4=μ√3/21/4=μ√3/2μ=1/√3/2